# KCL-TH-2024-59

### v1

@@ IV.1 Stationary boson stars
[…]
The remaining ground-state models represent a one-parameter family of BSs with each member determined by its central scalar amplitude $|\phi_{\rm c}|$. The solution space is conveniently displayed as a curve $M(|\phi_{\rm c}|)$ in the mass-amplitude222 Alternatively, the solutions can be represented in a mass-radius or $M(R)$ diagram. plane where, analogous to the Tolman-Oppenheimer-Volkoff Oppenheimer and Volkoff (1939); Tolman (1939) solutions for spherically symmetric NSs, it divides into stable and unstable branches Cook et al. (1994); Friedman et al. (1988); Harrison et al. (1965); Straumann (1984). These branches are delineated by local extrema of the function $M(|\phi_{\rm c}|)$ where the time dependence of radial peturbations changes from oscillatory to exponential Lee and Pang (1989); Gleiser (1988); Gleiser and Watkins (1989); Jetzer (1989a, 1990, b); Seidel and Suen (1990); Balakrishna et al. (1998); Alcubierre et al. (2003). For so-called mini BSs, where $\sigma\rightarrow\infty$, composed of a non-interacting scalar field, for example, the $M(|\phi_{\rm c}|)$ curve consists of one stable ($|\phi|<|\phi|_{\rm Kp}$) and one unstable branch ($|\phi|>|\phi|_{\rm Kp}$) separated by the maximum-mass or Kaup limit Kaup (1968) model with $M=M_{\rm Kp}\approx 0.633\,M_{\rm Pl}^{2}/m$ for $|\phi_{\rm c}|=|\phi|_{\rm Kp}\approx 0.0765\,M_{\rm Pl}$. For the solitonic potential (1), however, the branch structure becomes more complicated; more precisely, it mirrors the more complicated structure of the potential function $V(|\phi|)$ itself.
@@
Recalling from the above discussion of NSs and mini BSs that local extrema in the $M(|\phi_{\rm c}|)$ curve typically333 The rule is not rigorous; for example increasing $|\phi_{\rm c}|$ well beyond the Kaup limit for mini BSs ultimately leads to further extrema in $M(|\phi_{\rm c}|)$ for ultra-compact BSs, none of them marking a return to stable BSs. The presence or absence of a change in stability depends on whether the static perturbation belongs to the fundamental radial oscillation mode or an overtone Gleiser and Watkins (1989). mark the transition between stable and unstable models, stable BS models will be located on rising slopes of $M(|\phi_{\rm c}|)$, i.e. have $\mathrm{d}M/\mathrm{d}|\phi_{\rm c}|>0$ while unstable models mainly coincide with $\mathrm{d}M/\mathrm{d}|\phi_{\rm c}|<0$. Our crude analogy to homogeneous scalar fields would then suggest that minima (maxima) in $V(|\phi_{\rm c}|)$ coincide with rising (descending) slopes of $M(|\phi_{\rm c}|)$.
@@
Figure 4: $M(|\phi_{c}|)$ diagrams for the one-parameter families of BSs obtained for $\sigma/M_{\rm Pl}=0.2$, $0.25$, $0.28725$, $0.3$, $0.5$ as well as the mini-BS limit $\sigma\rightarrow\infty$. For each one-parameter family of BS models, globally (perturbatively) stable models are displayed in light (dark) copper color while unstable stars are marked in black. The blue dashed curves, quantified in units of $m^{2}M_{\rm Pl}^{2}$ on the right vertical axis, represent the respective potential functions (1). The vertical lines mark the extrema of the potential and the mass-amplitude curves.
@@ IV.2 Time evolutions of single boson stars
Let us first consider the case of mini BSs shown in the top row of the figure. In complete agreement with Fig. 4, the global maximum, i.e. Kaup limit, at $|\phi_{\rm c}|\approx 0.0765\,M_{\rm Pl}$ separates the stable and unstable branches: all models to the left of this point stay in their initial configuration (the blue symbols placed at migration time $t=0$) whereas all models to the right collapse to BHs (red symbols, marking UB instability). As expected, the migration time away from their initial state exhibits a pole at the Kaup limit, commensurate with the vanishing of the fundamental radial eigenmode’s frequency Gleiser and Watkins (1989), and rapidly drops with increasing compactness. In our simulations, models beyond the mass minimum at $|\phi_{\rm c}|\approx 0.33$, despite being on the potentially stable branch III of Fig. 5, ubiquitously collapse to BHs as shown by the time evolutions $|\phi|_{\rm max}(t)$ for selected BSs in the top right panel.
As we decrease $\sigma$ in the following rows of Fig. 6, we see the second bump in the $M(|\phi_{\rm c}|)$ curve moving towards smaller $|\phi_{\rm c}|$ and substantially increasing in magnitude. For all cases $\sigma\leq 0.3$, this second bump results in two sets of stable BSs – cf. the blue dots in the left panels of Fig. 6 – corresponding to branches I and III of Fig. 5. All models beyond the second bump, i.e. all models belonging to branch IV, are unstable and ubiquitously collapse to BHs (red dots).
@@ V.1 General behaviour
The first and perhaps most elementary of our observations in Fig. 1 is the enhanced radiative efficiency of BS collisions as compared to their BH counterparts. This has already been noticed for oscillatons in Ref. Helfer et al. (2019) (see their Fig. 1) and in Ref. Evstafyeva et al. (2023a) for unequal-mass BS collisions (see their Fig. 9). We interpret this feature in terms of two competing effects: the GW emission is expected to increase with (i) the binary constituents’ compactness and (ii) a higher degree of asymmetry around merger. BSs, being less compact but more vulnerable to deformation at merger than BHs, are thus expected to have their radiative efficiency reduced by the first effect but enhanced by the second. Empirically, our results confirm the observations of the above literature, namely that the net effect is optimized for moderate compactness considerably below that of BHs. In this section, we will for now assert this empirical result which we colloquially phrase as “mergers of deformable BSs are louder [^5] Obviously, when the BS are too squishy, they approach the weak field limit and GW production becomes less efficient.” in order to capture the fact that GW efficiency depends on both, compactness and the tendency of BSs to deform into highly asymmetrical configurations during merger. We will then return to conjecture on its origin in section V.2.
@@
Branch IIUB (red): Beyond point A, i.e. the first maximum in the $M(|\phi_{\rm c}|)$ curve of Fig. 5, we reach the red patch of the unstable branch II; our initial BSs are unstable, collapse to individual BHs during infall and form a BH merger remnant at coalescence according to the case 2 scenario of Fig. 7. The radiated energy function $E_{\rm GW}(|\phi_{\rm c}|)$ should consequently exhibit a discontinuous jump from a large BS value to the generic $E_{\rm GW}^{\rm BH}\approx 6\times 10^{-4}\,M_{\rm tot}$, as has already been observed for head-on collisions of oscillatons Helfer et al. (2019).
@@ V.2 GW Energy dependence on compactness
For question (a), we have argued in Sec. V.1 that in a BS merger a greater asymmetry of the coalescence phase generates stronger gravitational radiation. Observation (b) can then be explained by the argument that higher compactness implies lower deformability. What we mean by the vague term “deformability” is “the amplitude of the response of the BS system to gravitational and matter perturbations” as quantified through tidal deformability parameters; see e.g. Ref. Sennett et al. (2017). Since most of the GW emission in a head-on merger is generated during the coalescence phase, it follows that the larger the initial asymmetry of the apparent horizon or compact matter remnant, the higher the GW emission during the merger and ringdown would be. A high compactness, on the other hand, implies a deeper gravitational potential well where matter fields are less susceptible to perturbations during a merger – they are “stiffer”. Clearly, this effect is competing with the fact that systems need to be strongly gravitating before GWs can be produced since spacetime itself is also extremely stiff, hence there is a maximum – in the case of $\sigma/M_{\rm Pl}=0.28725$, this occurs at [^7] It has been argued in Ref. Okounkova (2020) that in the strong-field regime, non-linearities around merger are rapidly encompassed by a common horizon (see their Sec. II B) and thus will not reach a detector outside the event horizon. Our observation of enhanced GW energy from more asymmetric mergers is independent of this interpretation of captured non-linearities. $|\phi_{\mathrm{c}}|\approx 0.13$ where $E_{\rm GW}(|\phi_{\rm c}|)/M_{\rm tot}=0.0035$.
Finally, we can infer the point of maximal GW emission by recalling that $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ is extremal at transition point C connecting branches II and III, provided point C admits stable models in its neighbourhood. Ignoring for the moment the needles and troughs due to premature BH collapse in in $E_{\rm GW}(|\phi_{\mathrm{c}}|)$, we can furthermore argue as follows that the extremum is likely to be a maximum: First, we know that $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ is a rising function at very small $|\phi_{\mathrm{c}}|$. $E_{\rm GW}$ can then be minimal at point C only if it has a maximum somewhere between $|\phi_{\mathrm{c}}|=0$ and point C. To the left of point C, we would therefore already be in the regime where increasing compactness lowers GW emission. It is hard then to conceive of a physical mechanism that would lead to yet another reversal of $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ towards a rising slope. We cannot rigorously rule out that some such mechanism exists, but parsimony suggests the transition point C simply marks the transition from increasing $E_{\rm GW}$ (due to increasing compactness) towards decreasing $E_{\rm GW}$ (due to reduced asymmetry). Empirically, we find this view fully confirmed; if point C exists and admits stable BSs in its neighbourhood, it marks maximal GW emission.
@@ V.3 GW energy dependence on $V(\phi)$
Results for further values of the solitonic potential parameter $\sigma$ are presented in Ref. Ge (2024); all of these fall into one of the three regimes listed here.
The remarkably rish structure that we have encountered in the dependence of the GW energy on the BS parameters is not accompanied by a corresponding complexity in the GW signals. Rather, we universally observe rather standard head-on waveforms with a shape quite indistinguishable by eye from their BH counterparts. For example, we display in Fig. 9 the leading quadrupole mode of the Newman-Penrose scalar $\Psi_{4}$ for $\sigma=0.25\,M_{\rm Pl}$ and various scalar-field amplitudes including configurations of the needle regime. All signals have the overall structure, also seen for BHs, as for example in Fig. 8 of Ref. Sperhake (2007), and only differ significantly in their overall amplitude.
@@ V.4 Impact of the initial BS separation
In our discussion in section V.1, we have identified two features in the shape of the GW energy $E_{\rm GW}(|\phi_{\rm c}|)$ which we expect to vary with the initial separation of the binary. The first observation concerns BS mergers with correspondingly large GW emission on the early part of branch IIUB to the right of transition point A, despite the fact that they are unstable to BH collapse. How can this be? The answer lies in the inevitably finite infall time of our BS collisions; in Fig. 6, the left, third from top panel demonstrates that for $\sigma=0.28725$ and $|\phi_{\rm c}|\lesssim 0.06\,M_{\rm Pl}$, BSs are indeed unstable to BH collapse but take a long time to do so, longer, in fact, than the duration of their infall. They consequently merge as increasingly compact BSs but not yet as BHs. At $|\phi_{\rm c}|=0.068\,M_{\rm Pl}$ we finally reach BSs so unstable that they complete their individual collapse prior to encountering each other. This behaviour also explains the continuity of the curve $E_{\rm GW}(|\phi_{c}|)$ curve at $|\phi_{\rm c}|=0.068\,M_{\rm Pl}$ in Fig. 1; for each finite infall time, we can fine tune $|\phi_{\rm c}|$ to obtain any “desired” BS compactness at merger – note that these BSs are not in equilibrium and, hence, not constrained by the maximum compactness of equilibrium stars. We have tested this hypothesis by rerunning selected configurations in this regime with different initial separation $D$ but all other parameters unchanged. We indeed find in these simulations that the sharp drop in the GW energy towards $E_{\rm GW}^{\rm BH}$ moves towards larger $|\phi_{\mathrm{c}}|$ when we decrease $D$ whereas it converges to the transition point A in the limit $D\rightarrow\infty$.

### v2

@@ IV.1 Stationary boson stars
Many of the prominent features we observe in the dynamics and GW emission of BS head-on collisions find their origin in the properties of single BS spacetimes. We therefore start the discussion of our results by recalling the essential characteristics of spherically symmetric, stationary BS models. For a fixed scalar potential, Eq. (1) in our case, the space of spherically symmetric solutions is controlled by the central scalar-field amplitude $|\phi_{\rm c}|$. More specifically, for each value $|\phi_{\rm c}|$, there exists a countably infinite number of BS models characterized by the number $n$ of zero crossings of the radial profile $|\phi(r)|$: the ground-state BS with $n=0$ and the excited [^1] Not to be confused with “perturbed” or “dynamic” BSs which are stars evolving in time. Excited BSs, in contrast, are equilibrium solutions merely having one or more zero crossings in their radial scalar-field profile. stars for $n=1,\,2,\,\ldots$ . For a broad range of scalar potentials, the latter have been found to be intrinsically unstable to generic perturbations Balakrishna:1997ej, but more recent studies have found that sufficiently large self-interaction terms may cure such instabilities Sanchis-Gual:2021phr; Brito:2023fwr. In any case, excited BSs will play no further role in this paper.
The remaining ground-state models represent a one-parameter family of BSs with each member determined by its central scalar amplitude $|\phi_{\rm c}|$. The solution space is conveniently displayed as a curve $M(|\phi_{\rm c}|)$ in the mass-amplitude [^2] Alternatively, the solutions can be represented in a mass-radius or $M(R)$ diagram. plane where, analogous to the Tolman-Oppenheimer-Volkoff Oppenheimer:1939ne; Tolman:1939jz solutions for spherically symmetric NSs, it divides into stable and unstable branches Cook1994; Friedman1988; Harrison1965; Straumann1984. These branches are delineated by local extrema of the function $M(|\phi_{\rm c}|)$ where the time dependence of radial peturbations changes from oscillatory to exponential Lee:1988av; Gleiser:1988rq; Gleiser:1988ih; Jetzer:1988vr; Jetzer:1989qp; Jetzer:1989vs; Seidel:1990jh; Balakrishna:1997ej; Alcubierre:2003sx. For so-called mini BSs, where $\sigma\rightarrow\infty$, composed of a non-interacting scalar field, for example, the $M(|\phi_{\rm c}|)$ curve consists of one stable ($|\phi|<|\phi|_{\rm Kp}$) and one unstable branch ($|\phi|>|\phi|_{\rm Kp}$) separated by the maximum-mass or Kaup limit Kaup:1968zz model with $M=M_{\rm Kp}\approx 0.633\,M_{\rm Pl}^{2}/m$ for $|\phi_{\rm c}|=|\phi|_{\rm Kp}\approx 0.0765\,M_{\rm Pl}$. For the solitonic potential (1), however, the branch structure becomes more complicated; more precisely, it mirrors the more complicated structure of the potential function $V(|\phi|)$ itself.
@@
Recalling from the above discussion of NSs and mini BSs that local extrema in the $M(|\phi_{\rm c}|)$ curve typically [^3] The rule is not rigorous; for example increasing $|\phi_{\rm c}|$ well beyond the Kaup limit for mini BSs ultimately leads to further extrema in $M(|\phi_{\rm c}|)$ for ultra-compact BSs, none of them marking a return to stable BSs. The presence or absence of a change in stability depends on whether the static perturbation belongs to the fundamental radial oscillation mode or an overtone Gleiser:1988ih. mark the transition between stable and unstable models, stable BS models will be located on rising slopes of $M(|\phi_{\rm c}|)$, i.e. have $\mathrm{d}M/\mathrm{d}|\phi_{\rm c}|>0$ while unstable models mainly coincide with $\mathrm{d}M/\mathrm{d}|\phi_{\rm c}|<0$. Our crude analogy to homogeneous scalar fields would then suggest that minima (maxima) in $V(|\phi_{\rm c}|)$ coincide with rising (descending) slopes of $M(|\phi_{\rm c}|)$.
@@
Figure 4: $M(|\phi_{\rm c}|)$ diagrams for the one-parameter families of BSs obtained for $\sigma/M_{\rm Pl}=0.2$, $0.25$, $0.28725$, $0.3$, $0.5$ as well as the mini-BS limit $\sigma\rightarrow\infty$. For each one-parameter family of BS models, globally (perturbatively) stable models are displayed in light (dark) copper color while unstable stars are marked in black. The blue dashed curves, quantified in units of $m^{2}M_{\rm Pl}^{2}$ on the right vertical axis, represent the respective potential functions (1). The vertical dotted lines mark the extrema of the potential (blue) and the mass-amplitude curves (orange).
@@ IV.2 Time evolutions of single boson stars
Let us first consider the case of mini BSs shown in the top row of the figure. In complete agreement with Fig. 4, the global maximum, i.e. Kaup limit, at $|\phi_{\rm c}|\approx 0.0765\,M_{\rm Pl}$ separates the stable and unstable branches: all models to the left of this point stay in their initial configuration (the blue symbols placed at migration time $t=0$) whereas all models to the right collapse to BHs (red symbols, marking UB instability). As expected, the migration time away from their initial state exhibits a pole at the Kaup limit, commensurate with the vanishing of the fundamental radial eigenmode’s frequency [^4] We note Ref. Santos:2024vdm reports that extremal-mass configurations do not necessarily correspond to zero-frequency modes; our findings neither confirm nor contradict their results. Gleiser:1988ih, and rapidly drops with increasing compactness. In our simulations, models beyond the mass minimum at $|\phi_{\rm c}|\approx 0.33$, despite being on the potentially stable branch III of Fig. 5, ubiquitously collapse to BHs as shown by the time evolutions $|\phi|_{\rm max}(t)$ for selected BSs in the top right panel.
As we decrease $\sigma$ in the following rows of Fig. 6, we see the second bump in the $M(|\phi_{\rm c}|)$ curve moving towards smaller $|\phi_{\rm c}|$ and substantially increasing in magnitude. For all cases $\sigma\leq 0.3$, this second bump results in two sets of stable BSs – cf. the blue dots in the left panels of Fig. 6 – corresponding to branches I and III of Fig. 5. All models beyond the second bump, i.e. all models belonging to branch IV, are unstable and ubiquitously collapse to BHs (red dots); for the cases $\sigma\leq 0.3$ in Fig. 6, this transition to branch IV occurs at $|\phi_{\rm c}|\approx 0.22$.
@@ V.1 General behaviour
The first and perhaps most elementary of our observations in Fig. 1 is the enhanced radiative efficiency of BS collisions as compared to their BH counterparts. This has already been noticed for oscillatons in Ref. Helfer:2018vtq (see their Fig. 1) and in Ref. Evstafyeva:2022bpr for unequal-mass BS collisions (see their Fig. 9). We interpret this feature in terms of two competing effects: the GW emission is expected to increase with (i) the binary constituents’ compactness and (ii) a higher degree of asymmetry around merger. BSs, being less compact but more vulnerable to deformation at merger than BHs, are thus expected to have their radiative efficiency reduced by the first effect but enhanced by the second. Empirically, our results confirm the observations of the above literature, namely that the net effect is optimized for moderate compactness considerably below that of BHs. In this section, we will for now assert this empirical result which we colloquially phrase as “mergers of deformable BSs are louder [^6] Obviously, when the BS are too squishy, they approach the weak field limit and GW production becomes less efficient.” in order to capture the fact that GW efficiency depends on both, compactness and the tendency of BSs to deform into highly asymmetrical configurations during merger. We will then return to conjecture on its origin in section V.2.
@@
Branch IIUB (red): Beyond point A, i.e. the first maximum in the $M(|\phi_{\rm c}|)$ curve of Fig. 5, we reach the red patch of the unstable branch II; our initial BSs are unstable, collapse to individual BHs during infall and form a BH merger remnant at coalescence according to the case 2 scenario of Fig. 7. The radiated energy function $E_{\rm GW}(|\phi_{\rm c}|)$ should consequently exhibit a discontinuous jump from a large BS value to the generic $E_{\rm GW}^{\rm BH}\approx 6\times 10^{-4}\,M_{\rm tot}$, as has already been observed for head-on collisions of oscillatons Helfer:2018vtq.
@@ V.2 GW Energy dependence on compactness
For question (a), we have argued in Sec. V.1 that in a BS merger a greater asymmetry of the coalescence phase generates stronger gravitational radiation. Observation (b) can then be explained by the argument that higher compactness implies lower deformability. What we mean by the vague term “deformability” is “the amplitude of the response of the BS system to gravitational and matter perturbations” as quantified through tidal deformability parameters; see e.g. Ref. Sennett:2017etc. Since most of the GW emission in a head-on merger is generated during the coalescence phase, it follows that the larger the initial asymmetry of the apparent horizon or compact matter remnant, the higher the GW emission during the merger and ringdown would be. A high compactness, on the other hand, implies a deeper gravitational potential well where matter fields are less susceptible to perturbations during a merger – they are “stiffer”. Clearly, this effect is competing with the fact that systems need to be strongly gravitating before GWs can be produced since spacetime itself is also extremely stiff, hence there is a maximum – in the case of $\sigma/M_{\rm Pl}=0.28725$, this occurs at [^8] It has been argued in Ref. Okounkova:2020vwu that in the strong-field regime, non-linearities around merger are rapidly encompassed by a common horizon (see their Sec. II B) and thus will not reach a detector outside the event horizon. Our observation of enhanced GW energy from more asymmetric mergers is independent of this interpretation of captured non-linearities. $|\phi_{\mathrm{c}}|\approx 0.13$ where $E_{\rm GW}(|\phi_{\rm c}|)/M_{\rm tot}=0.0035$.
Finally, we can infer the point of maximal GW emission by recalling that $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ is extremal at transition point C connecting branches II and III, provided point C admits stable models in its neighbourhood. Ignoring for the moment the needles and troughs due to premature BH collapse in $E_{\rm GW}(|\phi_{\mathrm{c}}|)$, we can furthermore argue as follows that the extremum is likely to be a maximum: First, we know that $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ is a rising function at very small $|\phi_{\mathrm{c}}|$. $E_{\rm GW}$ can then be minimal at point C only if it has a maximum somewhere between $|\phi_{\mathrm{c}}|=0$ and point C. To the left of point C, we would therefore already be in the regime where increasing compactness lowers GW emission. It is hard then to conceive of a physical mechanism that would lead to yet another reversal of $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ towards a rising slope. We cannot rigorously rule out that some such mechanism exists, but parsimony suggests the transition point C simply marks the transition from increasing $E_{\rm GW}$ (due to increasing compactness) towards decreasing $E_{\rm GW}$ (due to reduced asymmetry). Empirically, we find this view fully confirmed; if point C exists and admits stable BSs in its neighbourhood, it marks maximal GW emission.
@@ V.3 GW energy dependence on $V(\phi)$
Results for further values of the solitonic potential parameter $\sigma$ are presented in Ref. Ge2024; all of these fall into one of the three regimes listed here.
The remarkably rich structure that we have encountered in the dependence of the GW energy on the BS parameters is not accompanied by a corresponding complexity in the GW signals. Rather, we universally observe rather standard head-on waveforms with a shape quite indistinguishable by eye from their BH counterparts. For example, we display in Fig. 9 the leading quadrupole mode of the Newman-Penrose scalar $\Psi_{4}$ for $\sigma=0.25\,M_{\rm Pl}$ and various scalar-field amplitudes including configurations of the needle regime. All signals have the overall structure, also seen for BHs, as for example in Fig. 8 of Ref. Sperhake:2006cy, and only differ significantly in their overall amplitude.
@@ V.4 Impact of the initial BS separation
In our discussion in section V.1, we have identified two features in the shape of the GW energy $E_{\rm GW}(|\phi_{\rm c}|)$ which we expect to vary with the initial separation of the binary. The first observation concerns BS mergers with correspondingly large GW emission on the early part of branch IIUB to the right of transition point A, despite the fact that they are unstable to BH collapse. How can this be? The answer lies in the inevitably finite infall time of our BS collisions; in Fig. 6, the left, third from top panel demonstrates that for $\sigma=0.28725$ and $|\phi_{\rm c}|\lesssim 0.06\,M_{\rm Pl}$, BSs are indeed unstable to BH collapse but take a long time to do so, longer, in fact, than the duration of their infall. They consequently merge as increasingly compact BSs but not yet as BHs. At $|\phi_{\rm c}|=0.068\,M_{\rm Pl}$ we finally reach BSs so unstable that they complete their individual collapse prior to encountering each other. This behaviour also explains the continuity of the curve $E_{\rm GW}(|\phi_{\rm c}|)$ curve at $|\phi_{\rm c}|=0.068\,M_{\rm Pl}$ in Fig. 1; for each finite infall time, we can fine tune $|\phi_{\rm c}|$ to obtain any “desired” BS compactness at merger – note that these BSs are not in equilibrium and, hence, not constrained by the maximum compactness of equilibrium stars. We have tested this hypothesis by rerunning selected configurations in this regime with different initial separation $D$ but all other parameters unchanged. We indeed find in these simulations that the sharp drop in the GW energy towards $E_{\rm GW}^{\rm BH}$ moves towards larger $|\phi_{\mathrm{c}}|$ when we decrease $D$ whereas it converges to the transition point A in the limit $D\rightarrow\infty$.